At the beginning of the first day (day 1) after grape harvesting is completed, a grape grower has 8000 kg of grapes in storage. On day n, for n = 1, 2, . . . , the grape grower sells 250n/(n + 1) kg of the grapes at the local market at the price of $2.50 per kg. He leaves the rest of the grapes in storage where each day they dry out a little so that their weight decreases by 3%. Let wn be the weight (in kg) of the stored grapes at the beginning of day n for n ≥ 1 (before he takes any to the market).




(a) Find the value of wn for n = 2.




(b) Find a recursive definition for wn. (You may find it helpful to draw a timeline.)




(c) Let rn be the total revenue (in dollars) earned from the stored grapes from the beginning of day 1 up to the beginning of day n for n ≥ 1. Find a recursive formula for rn.











1
Expert's answer
2022-06-07T14:37:35-0400

Solution:

Let's denote given values:

"w1" =800 kg - total grapes beginning of the 1st day (before sell);

"k1=(250*1)\/(1+1); k2=(250*2)\/(2+1); .....k(n-1)=(250*(n-1))\/((n-1)+1);"

and "kn=(250*n)\/(n+1);" weight of sold grapes 1st, 2nd ... and nth days (in kg).

"\alpha=0.03" - rate of weight decreases by drying.

a) If "wn" - weight of the stored grapes at the beginning of day nth, for n=2:

"w2=w1-k1-(w1-k1)*\alpha=(w1-k1)(1-\alpha);"

So,

"w2=(800kg-(250\/2)kg)*(1-0.03)=654.75kg."

b) We found for "w2=(w1-k1)(1-\alpha);"

so, for "w3=(w2-k2)(1-\alpha);"

if we continue this sequence, we can find:

"wn=(w(n-1)-k(n-1))(1-\alpha)."

c) We know:

"k1=(250*1)\/(1+1); k2=(250*2)\/(2+1); .....k(n-1)=(250*(n-1))\/((n-1)+1);"

and "kn=(250*n)\/(n+1);" weight of sold grapes 1st, 2nd ... and nth days (in kg).

If "rn" revenue up to beginning of the day n:

"rn=k1*2.5$+k2*2.5$+...+k(n-1)*2.5$;"

"rn=250*2.5$(1\/2+2\/3+3\/4+...+(n-1)\/n)."

Answer:

a) "w2=654.75 kg."

b) "wn=(w(n-1)-k(n-1))(1-\alpha)."

c) "rn=(1\/2+2\/3+3\/4+...+(n-1)\/n)*625$."